Circular No. 7826

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 B3 (LINEAR) F. Shelly, Lincoln Laboratory, Massachusetts Institute of Technology,

reports that an apparently asteroidal object discovered by LINEAR on Jan. 26 (and placed on the NEO Confirmation Page) was found to show a clear tail in p.a. ∼ 330^{o} on Feb. 11.09 UT. T. B. Spahr, Harvard-Smithsonian Center for Astrophysics, reports that CCD images taken with the 1.2-m reflector at Mt. Hopkins on Feb. 12.13 show the object to be very faintly diffuse with a faint extension in p.a. 315^{o}.

2002 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 26.12202  2^{h}49^{m}14.^{s}23  +79^{o}11′59′′.1  19.5

The available astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2002-C81:

T = 2002 Jan. 31.518 TT  ω = 124.^{o}493
Ω = 289.526 }2000.0
q = 6.04878 AU  i = 73.590

SUPERNOVA 2002ap IN M74 Y. Yoshii, T. Minezaki, T. Aoki, K. Enya, M. Doi, and K. Motohara,

Institute of Astronomy, University of Tokyo; Y. Kobayashi, National Astronomical Observatory of Japan; M. Suganuma and H. Tomita, Department of Astronomy, University of Tokyo; and B. A. Peterson, Australian National University, report the following magnitudes (uncertainty 0.1–0.2 mag, optical magnitudes on Johnson-Cousins system) obtained with an optical and near-infrared camera (MIP) at the Cassegrain focus of the MAGNUM 2m telescope at Haleakala Observatory, Maui: Feb. 2.408 UT, B = 13.34; 2.420, J = 12.54; 2.431, V = 12.82; 2.438, R = 12.79; 2.445, K = 12.31; 2.450, U = 13.51; 4.396, B = 13.15; 4.400, J = 12.27; 4.405, V = 12.54; 4.413, U = 13.23; 4.417, K = 11.94; 4.422, R = 12.51.

SUPERNOVA 2002H IN MCG –02-35-11 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002H (cf. IAUC 7798, 7815), obtained by M. Calkins on Jan. 22.54 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova, ∼ three weeks past maximum.

2002 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green